Solve the system of equations.
4x + y = 11
16x + 4y = 44

Answer by checkley77(12844)

4x + y = 11 multiply this equation by -4 & add
16x + 4y = 44
-16x-4y=-44
----------------
0x+0x=0 no unique solution as thes etwo are identical equations with 2 unknowns.